# A Survey on Integration of Large Language Models with Intelligent Robots

### v2

@@ 4.2 Sensing Behavior
On the other hand, active perception refers to the conscious process of gathering sensory information by taking additional actions. Active information gathering enhances environmental understanding by acquiring new information through sensory observations or requesting user feedback [129, 79]. For example, LLM-Planner [129] generates seeking actions such as ‘open the refrigerator’ to locate invisible objects. Recent studies also focus on collecting sensory data to better understand objects’ physical properties [168, 48]. However, LLMs often generate inaccurate or fabricated information, known as hallucinations. To address this issue, Dai et al. [26] introduce a personalized conversational agent designed to ask users for uncertain information.
@@ 5.1 Task Planning
LLM-based task planners are capable of generating plans without strict symbol definitions [58], while traditional task planners are required to pre-define operators with domain knowledge about available actions and constraints [99, 35]. In this field, most planners employ a static planning strategy, which takes fixed descriptions that are not adaptable to changes in environment [163]. However, an alternative approach, adaptive planning, allows for the incorporation of environmental feedback into input prompts, enabling adjustments to actions based on observed conditions. This section reviews LLM-based planners in terms of these two strategies: static and adaptive planning.
Static planning: Static planning approaches are generally zero- or few-shot prediction methods, where zero-shot methods generate a plan based solely on an input command, while few-shot methods leverage learning from a limited set of similar examples [10, 28, 70, 163]. However, LLMs often exhibit poor performance in long-horizon task planning due to limited reasoning ability [90, 139]. To address this limitation, Huang et al. [58] introduce a planner that iteratively selects the most probable action among executable ones generated by LLMs.
Alternatively, LLM-based code generators, such as Code as Policies [87] or ProgPrompt [128], produce codes that result in adaptive actions responsive to observations [57, 56]. Singh et al. [128] demontrate that code generation outperforms basic task planning from LLMs since the output plan closely aligns with execution environments. Despite their advantages, these methods lack validation and replanning processes.
To validate plans, researchers often augment LLMs with logical programs, either to (1) check if resulting plans violate logical constraints or (2) generate plans using an external logical planner. For instance, SayPlan [118], a GPT4-based planner, validates abstract-level actions through a scene-graph simulator 3DSG [1], while LLM+P [90] applies a PDDL problem translated from LLMs to a classical task planner, Fast Downward [45]. In addition, Silver et al. [127] demonstrate that a search-based planner with an initial plan from LLMs performs better by exploring fewer nodes. These studies underscore the effectiveness of integrating LLMs with logical programs to increase the success rate or the performance of generating feasible plans.
Adaptive planning: Adaptive planning allows robots to modify their plans or actions in response to feedback, either by generating new plans based on environmental observations [21, 151, 141, 168, 169] or by detecting failures and adjusting accordingly [61]. Chen et al. [13] and Huang et al. [60] introduce adaptation strategies that generate new plans based on observed feedback, enabling robots to respond to a broader range of scenarios.
@@ 5.2 Task and Motion Planning
Motion planning refers to generating an objective of trajectory with numerical waypoints within a robot’s configuration space or task space. However, direct numerical sequencing is challenging since the language models learn through generating tokens unrelated to continuous space. Despite this, an LLM-based motion planner directly generates positional sequences for drone choreography [68], as their task is simple enough to demonstrate LLMs’ spatial reasoning ability. For more complex scenarios with an indirect approach, Huang et al. [59] augment LLMs with a search-based planner. In their framework, VoxPoser, an LLM generates code of potential field using VLM, and then the search-based planner conducts motion planning within the generated field.
TAMP refers to integrating high-level task planning with low-level motion planning. Various works use LLMs themselves as TAMP planners, exploiting both logical and physical reasoning capabilities [97, 80, 152]. Researchers guide LLMs to generate high-level subgoals and then use them for low-level trajectory generation [97, 80]. However, their coarse representations limit their methods to simple tasks such as pick-and-place. Instead, Xia et al. [152] enhance LLMs’ kinematic knowledge using kinematic-aware prompting for complex manipulation such as articulated object manipulation. In addition, various studies augment LLMs to complement their reasoning abilities. Researchers often integrate a logic-augmented TAMP planner that checks the logical feasibility of the task plan [30]. Meanwhile, others use physics-augmented TAMP planner to evaluate physical feasibility [89, 19, 44]. Text2Motion [89], for example, allows an LLM to generate physically-feasible high-level actions and combine them with learned skills for low-level actions.
@@ 6 Control
Early studies primarily focused on establishing mappings between simple linguistic commands and known motion primitives. With the advent of deep learning, researchers have explored two main approaches in control: direct modeling of control values based on linguistic instructions [119, 8] and indirect interpretation of complex instructions via LLMs to generate actions [153]. We categorize the work in this field into two groups: (1) direct approach which means the direct generation of control commands based on linguistic instructions and (2) indirect approach which stands for indirect specification of control commands through linguistic guidance. Fig. 1 presents a detailed categorization alongside related papers, referred in orange cells.
@@ 6.1 Direct Approach
The direct approach involves using an LLM to interpret and produce executable commands, either by selecting motion primitives [133] or generating control signals [170, 145]. Early work generates action tokens to produce a control policy through training Transformer architecture [140] with task-specific expert demonstration data [119, 8, 131]. Researchers linearly map these tokens to discretized end-effector velocities [119] or displacements [8, 131] for continuous motion. While these approaches demonstrate a degree of generalization over unseen tasks, such as new objects or realistic instructions, they often require extensive data collection and training time.
To reduce the collection effort, researchers often leverage existing web-scale vision and language datasets for finetuning [170, 142]. For example, Zitkovich et al. co-finetune VLMs, such as PaLI-X [18] and PaLM-E [32], targeting both visual-language tasks and robotics control tasks. They use the original datasets designed for VLMs alongside robotics control demonstrations to preserve general knowledge during finetuning, as demonstrated in RT-2 [170]. Additionally, to reduce the training burden, Hu et al. [51] use a low-rank adaptation (LoRA) method for finetuning an LLM for control tasks [16] rather than finetuning the entire model.
LLMs often struggle to generate continuous action-level commands such as joint position and torque values, as LLMs typically generate atomic elements known as tokens [133]. Therefore, researchers instead generate task-level output using LLMs [133, 102, 11]. For example, SayTap, an LLM-based controller for walking, generates contact patterns between feet and the ground for walking motion with an LLM instead of directly producing joint positions [133]. Other studies address the control problem by framing it as completing a sequence of end-effector poses [102] or generating Python code [11], similar to natural-language generation tasks. Recently, researchers prompt an LLM to produce action space output by providing normalized and discretized control value history to maintain continuity in control [145] or by providing robot kinematics information to determine reasonable joint values for a desired pose [86].
@@
Table 1: A conversational prompt for interactive grounding. Through the “Task” in the prompt, we ask the LLM to ground the underspecified object, referred to as “something” in the “Task”, as a “cookie” by interactively asking for personal preferences. The prompt consists of task description, task procedure, and task context parts guiding the LLM’s behavior and contextual understanding. The words in bold indicate subjects of interactions with LLM responses highlighted in blue.
@@ 6.2 Indirect Approach
LLMs are also useful for generating indirect representations of control commands (e.g., subgoals or reward functions), based on natural language instructions. Researchers leverage goal descriptions, explaining desired behaviors in natural language, to guide the learning process [33, 67, 78]. For example, ELLM [33], an LLM-based RL framework, uses an LLM to generate subgoal descriptions as a condition of the RL policy and further uses the similarity between current observation and the subgoal description in text embedding space to calculate reward. Further, Kumar et al. [78] incrementally use an LLM to generate a goal description based on previous human instructions. However, as the output of an LLM is a natural-language description, these approaches require an additional step of grounding or interpreting the description.
Using code generation capability, researchers generate a code-level reward function, instead. Yu et al. [161] convert a natural-language goal into high-level motion descriptions and then generate a corresponding reward function. However, this generation requires fixed reward-function formats. Instead, recent work prompts an LLM to infer reward-function formats from human-designed examples [71, 144]. Nonetheless, the generated reward functions may not always be accurate or optimal enough for direct use in training [130].
To improve accuracy, researchers add a refinement loop to validate both the syntax [112] and semantics [153, 130, 96, 165] of the generated reward functions. For example, Song et al. [130] use an LLM to re-design a reward function based on the convergence of the training process and resulting robot motion. Further, researchers use an LLM to evaluate robot motion, directly generating rewards [25]. In addition, recent work demonstrates the use of an LLM in refining motion by adjusting control parameters based on the error state [132] or by selecting a suitable motion target from human feedback [91].
@@ 7 Prompt Guideline
We provide prompt design guidelines for robotic tasks to researchers entering this field. A prompt is a message to guide LLMs to process and output as we instruct pre-trained language models [93, 149]. Well-designed prompts
@@ 7.1 Conversation Prompt: Interactive Grounding
We detail a conversation prompt design, leveraging an LLM as a grounding agent, to clarify commands such as “Bring me something to eat” and infer the ambiguous targets, expressed as “something,” through logical inference. Table 1 shows the design detail, where the prompt consists of three key components: task description, task procedure, and task context. We further describe each as follows.
@@
The task context describes the contextual inputs, such as “world model,” that LLMs perform grounding upon. Consistency in terminology across task description and task procedure is crucial for LLM operations. For example, the common expressions, such as “tasks” and “world model” allow the LLM to work within the same context provided. Further, by using clear names for objects in the world model, we enable the LLM to apply common knowledge to named entities. Note that, although we use a list of objects as a world model, LLMs accept world models in various formats: textual descriptions, a list of objects, and scene graphs.
With these structured components, the prompt invokes an interactive grounding dialogue for precise object identification, as shown in Table 1. This prompt utilizes ChatGPT 3.5 [106].
@@
You are an agent responsible for inferring edge relations in a scene graph. Based on the provided names and 3D coordinate information of each object, infer one major {spatial_relation} from {source} to {target} (i.e., source->>>target: spatial_relation), which means source is located to spatial_relation of target. Spatial relations are limited to 'left', 'right', 'forward', 'back', 'up', and 'down'. The bbox_extent represents the object's dimensions along the X, Y, and Z axes in meters, and the bbox_center specifies the object's central position in 3D space. For example, if the y-value of the source's center is bigger than that of the target, then the source is located to the forward of the target (i.e., source->>>target: forward). Similarly, if the x-value of source is bigger than the target’s, then the source is located to the right of the target (i.e., target->>>source: right). You can determine 'left' and 'right' based on the x-value, 'forward' and 'back' based on the y-value, and 'up', 'down' based on the z-value.
@@
Provide the output in a simple format as shown in the example below. Take a deep breath and work on this problem step-by-step.
@@
LLM Output Scene Graph Visualization
@@
Table 2: Directive Prompts for generating a scene graph. The table includes two prompts: node creation and edge creation. Given images in a scene, a multimodal LLM perceives objects and infers relevant relationships using geometric information. The words in bold indicate subjects of outputs with LLM responses highlighted in blue. The visualizations are neither the inputs nor the outputs of the LLMs.
7.2 Directive Prompt: Scene Graph Generation
We introduce directive prompt designs for constructing a scene graph from a scene image using a multimodal LLM, particularly with GPT-4 [107]. The scene graph consists of objects as nodes and their relationships as edges. Despite the advancement of multimodal LLMs, their capability has limitations in inferring 3D relationships from a 2D image [14]. To reduce the limitation, we decompose the task into two steps: node creation with multimodal inputs and edge creation with textual information. We describe each step with detailed examples provided in Table 2.
@@ 7.2 Directive Prompt: Scene Graph Generation
Create a next plan for a robot to complete a household task using the allowed actions and visible objects. The robot is able to pick up and hold only one object. The robot CANNOT open or close some objects when the robot holds an object (after PickUp and before PutInto). Initially, the robot does not hold anything and every openable object is closed. All plan sequence starts with 'Initial' which means initial condition. When the task is completed, use the action 'Done' to finish it. Do not do other than the task requires. Answer only one step of next plan without other things.
@@
The prompt for node creation consists of two parts: (1) task description and (2) task context. The task description includes the LLM’s expected behavior (i.e., role) and response format as Section 7.1. For instance, the LLM’s role is to identify objects as nodes in the given image. We then specify the output format as ‘ObjectName(ID)’ for consistency and simplicity. Then, the task context presents a sequence of unique object identifiers with corresponding object-centric images. We obtain the object-centric images from a scene by cropping them via SAM [77], a foundation vision model proficient in identifying objects with occlusion.
The edge creation consists of (1) task description, (2) examples, and (3) task context. The task description not only specifies the expected behavior and output format but also elucidates on how to identify the relationships between nodes leveraging examples. We particularly explain how the LLM uses 3D object coordinates and unit measurements to infer spatial relationships from a predefined set such as ‘left,’ ‘right,’ etc. Unlike the node creation, this allows for the generation of additional output explanations to accommodate the complexity discerning spatial relationships.
To enhance the understanding of the input format and corresponding output, we include examples showcasing edge generation. We choose an example similar to the target scenario in terms of objects and their spatial interrelationships, thereby providing richer information for edge identification. Finally, the task context provides source and target node information as inputs and empty output to obtain responses from the LLM. Instead of providing a whole permutation of nodes, we specify a whole combination of nodes, assuming that we can identify the edge of the reversed direction as the opposite spatial relationship (e.g., ‘left’ is opposite to ‘right’).

### v3

@@ 4.2 Sensing Behavior
On the other hand, active perception refers to the conscious process of gathering sensory information by taking additional actions. Active information gathering enhances environmental understanding by acquiring new information through sensory observations or requesting user feedback [129, 78]. For example, LLM-Planner [129] generates seeking actions such as ‘open the refrigerator’ to locate invisible objects. Recent studies also focus on collecting sensory data to better understand objects’ physical properties [168, 48, 132]. However, LLMs often generate inaccurate or fabricated information, known as hallucinations. To address this issue, Dai et al. introduce a personalized conversational agent designed to ask users for uncertain information [25].
@@ 5.1 Task Planning
LLM-based task planners are capable of generating plans without strict symbol definitions [58], while traditional task planners are required to pre-define operators with domain knowledge about available actions and constraints [98, 34]. In this field, most planners employ a static planning strategy, which takes fixed descriptions that are not adaptable to the changes in environment [163]. However, an alternative approach, adaptive planning, allows for the incorporation of environmental feedback into input prompts, enabling adjustments to actions based on observed conditions. This section reviews LLM-based planners in terms of these two strategies: static and adaptive planning.
Static planning: Static planning approaches are generally zero- or few-shot prediction methods, where zero-shot methods generate a plan based solely on an input command, while few-shot methods leverage learning from a limited set of similar examples [9, 27, 70, 163]. However, LLMs often exhibit poor performance in long-horizon task planning due to limited reasoning ability [89, 140]. To address this limitation, Huang et al. introduce a planner that iteratively selects the most probable action among executable ones generated by LLMs [58].
Alternatively, LLM-based code generators, such as Code as Policies [86] or ProgPrompt [128], produce codes that result in actions responsive to observations [57, 56]. Singh et al. demonstrate that code generation outperforms basic task planning from LLMs since the output plan closely aligns with execution environments [128]. Despite their advantages, these methods lack validation and replanning processes.
To validate plans, researchers often augment LLMs with logical programs, either to (1) check if resulting plans violate logical constraints or (2) generate plans using an external logical planner. For instance, SayPlan [118], a GPT4-based planner, validates abstract-level actions through a scene-graph simulator 3DSG [1], while LLM+P [89] applies a PDDL problem translated from LLMs to a classical task planner, Fast Downward [45]. In addition, Silver et al. demonstrate that a search-based planner with an initial plan from LLMs performs better by exploring fewer nodes [127]. These studies underscore the effectiveness of integrating LLMs with logical programs to increase the success rate or the performance of generating feasible plans.
Adaptive planning: Adaptive planning allows robots to modify their plans or actions in response to feedback, either by generating new plans based on environmental observations [20, 152, 142, 168, 169] or by detecting failures and adjusting accordingly [61]. Chen et al. and Huang et al. introduce adaptation strategies that generate new plans based on observed feedback, enabling robots to respond to a broader range of scenarios [12, 60].
@@ 5.2 Task and Motion Planning
Motion planning refers to the process of generating a path by computing sequential waypoints within configuration or task spaces. Jiao et al. introduce an LLM-based motion planner that directly generates positional sequences for drone choreography [68]. While this work demonstrates LLMs’ spatial reasoning abilities, the scenario presented is relatively simple. Further, planning spaces are often continuous, which presents a challenge for language models that operate with discrete tokens. Alternatively, indirect sequencing approaches, such as VoxPoser [59], generate a potential field code with the help of a VLM and then conduct motion planning within the generated field, augmenting LLMs with a search-based planner.
TAMP refers to integrating high-level task planning with low-level motion planning. Recent studies often use LLMs as TAMP planners, leveraging LLMs’ logical and physical reasoning capabilities [96, 79, 153]. Researchers guide LLMs to generate high-level subgoals, which are then used for low-level trajectory generation [96, 79]. However, the coarse representations of LLMs restrict their applications to simple tasks such as pick-and-place. To address this limitation, researchers use additional prompts or augment LLMs to improve reasoning abilities. For example, Xia et al. enable LLMs to consider kinematic knowledge through kinematic-aware prompting for more complex manipulation tasks, such as articulated object manipulation [153]. Ding et al. introduce a logic-augmented LLM planner that checks the logical feasibility of the task plans generated by LLMs [29]. Meanwhile, others use physics-augmented LLM planners to evaluate physical feasibility [88, 18, 44]. For example, Text2Motion [88] allows an LLM to generate physically feasible high-level actions and combine them with learned skills for low-level actions.
@@ 6 Control
Early studies primarily focus on establishing mappings between simple linguistic commands and known motion primitives. With the advent of deep learning, researchers have explored two main approaches in control: direct modeling of control values based on linguistic instructions [119, 7] and indirect interpretation of complex instructions via LLMs to generate actions [154]. We categorize the work in this field into two groups: (1) direct approach which means the direct generation of control commands based on linguistic instructions and (2) indirect approach which stands for indirect specification of control commands through linguistic guidance. Fig. 1 presents a detailed categorization alongside related papers, referred in orange cells.
@@ 6.1 Direct Approach
The direct approach involves using an LLM to interpret and produce executable commands, either by selecting motion primitives [134] or generating control signals [170, 146]. Early work generates action tokens to produce a control policy training Transformer architecture with task-specific expert demonstrations such as Gato [119], RT-1 [7], and MOO [131]. Researchers linearly map action tokens to discretized end-effector velocities [119] or displacements [7, 131] for continuous motion. While these approaches demonstrate a degree of generalization over unseen tasks, such as new objects or realistic instructions, they often require extensive data collection and training time.
To reduce the collection effort, researchers often leverage existing web-scale vision and language datasets such as RT-2 [170] and RT-X [143]. For example, Zitkovich et al. train VLMs (e.g., PaLI-X [17] and PaLM-E [31]) with visual-language datasets and robotic demonstrations [170]. This approach maintains general knowledge of visual-language tasks while training for control tasks. In addition, to reduce the training burden, Chen et al. use a low-rank adaptation (LoRA) [51] method for fine-tuning an LLM for control tasks rather than fine-tuning the entire model [15].
LLMs often struggle to generate continuous action-level commands such as joint position and torque values, as LLMs typically generate atomic elements known as tokens [134]. Therefore, researchers instead generate task-level output using LLMs [134, 101, 10]. For example, SayTap, an LLM-based controller for walking, generates contact patterns between feet and the ground for walking motion with an LLM instead of directly producing joint positions [134]. Other studies solve the control problem as natural-language generation tasks, completing a sequence of end-effector poses [101] or generating Python codes [10]. Recent studies often restrict action space to enhance LLM control outputs. For instance, Wang et al. design a prompt that produces positive integer control values while maintaining a smoothness trend in outputs [146]. Alternatively, Li et al. demonstrate that incorporating robot kinematics information helps the LLM determine joint values for desired poses [85].
@@
Table 1: A conversational prompt for interactive grounding. Through the ‘task’ in the prompt, we ask an LLM to ground the underspecified object, referred to as ‘something’ in the ‘task’ as a ‘cookie’ by interactively asking for a personal preference.
The prompt consists of task description, task procedure, and task context parts guiding the LLM’s behavior and contextual understanding. The words in bold indicate subjects of interactions with LLM responses highlighted in blue.
@@ 6.2 Indirect Approach
LLMs are also useful for generating indirect representations of control commands, such as subgoals or reward functions, based on natural language instructions. To guide the learning process, researchers leverage the goal description that explains desired behaviors in natural language [32, 67, 77]. For example, ELLM [32], an LLM-based reinforcement learning (RL) framework, uses an LLM to generate subgoal descriptions as prior knowledge of the RL policy and further uses the similarity between current observation and the subgoal description in text embedding space to calculate reward. Further, Kumar et al. generate a goal description based on the history of human instructions to reuse previously learned skills [77]. However, as the output of an LLM is a natural-language description, these approaches require an additional step of grounding or interpreting the description.
Alternatively, researchers often generate code-level reward functions. Yu et al. convert a natural-language goal into a high-level motion description and then generate a corresponding reward function [161]. However, this method requires pre-defined reward formats. Instead, recent work prompts an LLM to infer a new reward function from human-designed examples [71, 145]. Nonetheless, the generated reward functions may not always be accurate or optimal enough for direct use in training [130].
To improve accuracy, researchers add a refinement loop to validate both the syntax [112] and semantics [154, 130, 95, 165] of the generated reward functions. For example, Song et al. use an LLM to redesign a reward function based on the convergence of the training process and the resulting robot motion [130]. Chu et al. employ an LLM to directly generate rewards for evaluating robot motion [24]. Other approaches refine a motion by adjusting control parameters based on the error state [133] or by selecting a suitable motion target from human feedback [90].
@@ 7 Prompt Guideline
We provide prompt design guidelines for robotic tasks to researchers entering this field. A prompt is a message that directs LLMs to process and generate outputs according to our instructions [92, 150]. Well-designed prompts
@@ 7.1 Conversation Prompt: Interactive Grounding
We detail a conversation prompt design, leveraging an LLM as a grounding agent, to clarify commands such as ‘Bring me something to eat’ and infer the ambiguous targets, expressed as ‘something,’ through logical inference. Table 1 shows the design detail, where the prompt consists of three key components: task description, task procedure, and task context. We further describe each component as follows.
@@
The task context describes the contextual inputs, such as ‘world model,’ that LLMs perform grounding upon. Consistency in terminology across task description and task procedure is crucial for LLM operations. For example, the common expressions, such as ‘task’ and ‘world model’ allow the LLM to work within the same context provided. Further, by using clear names for objects in the world model, we enable the LLM to apply common knowledge to named entities. Note that, although we use a list of objects as a world model, LLMs accept world models in various formats: textual descriptions, a list of objects, and scene graphs.
With these structured components, the prompt invokes an interactive grounding dialogue for precise object identification, as shown in Table 1. We obtain the resulting interaction using ChatGPT 3.5 [106].
@@
You are an agent responsible for inferring edge relations in a scene graph. Based on the provided names and 3D coordinate information of each object, infer one major {spatial_relation} from {source} to {target} (i.e., source->>>target: spatial_relation), which means the source is located to spatial_relation of the target. Spatial relations are limited to 'left,' 'right,' 'forward,' 'back,' 'up,' and 'down'. The bbox_extent represents the object's dimensions along the X, Y, and Z axes in meters, and the bbox_center specifies the object's central position in 3D space. For example, if the y-value of the source's center is bigger than that of the target, then the source is located to the forward of the target (i.e., source->>>target: forward). Similarly, if the x-value of the source is bigger than the x-value of the target, then the source is located to the right of the target (i.e., target->>>source: right). You can determine 'left' and 'right' based on the x-value, 'forward' and 'back' based on the y-value, and 'up' and 'down' based on the z-value.
Provide the output in a simple format, as shown in the example below. Take a deep breath and work on this problem step-by-step.
@@
LLM Output Scene-Graph Visualization
@@
Table 2: Directive Prompts for generating a scene graph. The table includes two prompts: node creation and edge creation. Given segmented images in a scene, a VLM identifies objects, and then an LLM infers relevant relationships using geometric information. The words in bold indicate subjects of outputs with the VLM or LLM responses highlighted in blue. The visualizations are not part of any conversations.
7.2 Directive Prompt: Scene-Graph Generation
We introduce directive prompt designs for constructing a scene graph from a scene image using a multimodal LLM, particularly with GPT-4 [107]. The scene graph consists of objects as nodes and their relationships as edges [36, 41]. Despite the advancement of multimodal LLMs, their capability has limitations in inferring 3D relationships from a 2D image [13]. To reduce the limitation, we decompose the task into two steps: node creation with multimodal inputs and edge creation with textual information. We describe each step with detailed examples provided in Table 2.
@@ 7.2 Directive Prompt: Scene-Graph Generation
Create a next plan for a robot to complete a household task using the allowed actions and visible objects. The robot is able to pick up and hold only one object. The robot CANNOT open or close some objects when the robot holds an object (after PickUp and before PutInto). Initially, the robot does not hold anything, and every openable object is closed. All plan sequences start with 'Initial,' which means initial condition. When the task is completed, use the action 'Done' to finish it. Do not do anything other than what the task requires. Answer only one step of the next plan without other things.
@@
The prompt for node creation consists of two parts: (1) task description and (2) task context. The task description includes the multimodal LLM’s expected behavior (i.e., role) and response format, similar to Section 7.1. For instance, the multimodal LLM’s role is to identify objects as nodes in the given images. We then specify the output format as ‘ObjectName(ID)’ for consistency and simplicity. Then, the task context presents a sequence of unique object identifiers with corresponding object-centric images. In this scenario, we assume that object-centric images are obtained using an active perception method to identify objects with occlusion.
The edge creation consists of (1) task description, (2) examples, and (3) task context. The task description not only specifies the expected behavior and output format but also elucidates on how to identify the relationships between nodes leveraging examples. We particularly explain how the LLM uses 3D object coordinates and unit measurements to infer spatial relationships from a predefined set such as ‘left,’ ‘right,’ etc. Unlike the node creation, this allows for generating additional output explanations to accommodate the complexity of discerning spatial relationships.
To enhance the understanding of the input format and corresponding output, we include examples showcasing edge generation. We choose an example similar to the target scenario in terms of objects and their spatial interrelationships, thereby providing richer information for edge identification. Finally, the task context provides source and target node information as inputs and leaves the output empty to obtain responses from the LLM. We also assume that the 3D bounding boxes are obtained through a neural network-based detector [104] or from the extent of point clouds using depth information [41].